1. The sum of three numbers is 70. Two thirds of the second number is the first number. The last number is three times the second. Find the three numbers
2. the sum of three numbers is 276. The second number is one more than the first number, and the third number is two more than the first number. Find the three numbers.
3. there are 17 coins. (all nickles, dimes and quarters). The total vale of the coins is $2.40. Twice the number of nickles plus three times the number of dimes decreased by the number of quarters is 23. Find the number of each coin.

Answer by macston(5194)

.
1.)
S=second number; F=first number=(2/3)S; T=third number=3S
.
F+S+T=70
(2/3)S+S+3S-70
(14/3)S=70
S=(3/14)70=15
ANSWER 1: The second number is 15.
.
F=(2/3)S=(2/3)15=10
ANSWER 2: The first number is 10.
.
T=3S=3(15)=45
ANSWER: The third number is 45.
.
CHECK:
F+S+T=70
10+15+45=70
70=70
.
IN THE FUTURE, PLEASE LIMIT TO ONE QUESTION PER POST.
2.)
.
F=first number; S=second number=F+1; T=third number=S+2
.
F+S+T=276
F+(F+1)+(F+2)=276
3F+3=276
3F=273
F=91
ANSWER 1: The first number is 91
.
S=F+1=91+1=92
ANSWER 2: The second number is 92.
.
T=F+2=91+2=93
ANSWER 3: The third number is 93.
.
CHECK:
F+S+T=276
91+92+83=276
276=276
.
IN THE FUTURE, PLEASE LIMIT TO ONE QUESTION PER POST.